• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name kelleher

    Log entry time 09:34:58 on April 15, 2011

    Entry number 355591

    keyword=Owl Shift+ Summary

    Saturday, March 12, 2011
    
    Plan: Production data for SRC
    
    00:00 Shift begins in controlled access to move right spectrometer to 20
    degrees.
    00:20 Everyone out of the hall, restoring beam
    00:43 Turned off RHRS dipole regulation.  P0= 1.09908, I = 376.89
    03:43 Rebooted TS11 and TS0.  The were recording zero events, which may
    be normal for configuration UB2ArmNTS11.  After reboot, still no TS11 or TS0
    08:00 Collect optics data before turning off beam for pass change
    
    Run Summary
    
    Run   Begin  End    Notes
    -----------------------------------
    3530  00:43  00:46  spot++
    3531  00:51  00:55  rate check (2uA and 10uA)
    3532  00:56  01:19  peak search (10uA)
    3533  01:23  02:08  Production data
    3534  02:08  02:54  Production data  (
    3535  02:55  03:40  Production data (10uA)
    3536  03:43  04:28  Production data (10uA)
    3537  04:28  05:17  Production data (10uA)
    3538  05:18  06:03  Production data (10uA)
    3539  06:03  06:49  Production data (10uA)
    3540  06:50  07:16  Production data (10uA)
    3541  07:16  07:19  Rate check (40uA)
    3542  07:19  07:26  Rate Check (40uA)
    3543  07:26  07:55  High Current Run
    3544  07:55  08:05  BeO optics check prescales(BB on, raster on)
    3545  08:05  08:19  BeO optics (BB on, raster on)
    3546  08:22  08:31  BeO optics (BB on, raster off)
    3547  08:31  08:51  BeO optics (BB off, raster off)
    3548  08:54  09:05  BeO optics (BB off, raster on)
    3549  09:10  09:11  Multifoil optics (BB off, raster off)
    3550  09:11  09:22  Multifoil optics (BB off, raster off)
    3351  09:22  09:32  Multifoil optics (BB on, raster off)
    
    That's All, Folks!